We are using a Moving Average Forecast with n=4 and know that F10 = 49, D10 = 47, and F11 = 54. Find D6.
Answer #1

Answer: 27

Explanation:

F10 = (D6 + D7 + D8 + D9)/4

D6 + D7 + D8 + D9 = 4*49 = 196

F11 = (D7 + D8 + D9 + D10)/4

D7 + D8 + D9 + D10 = 4*54 = 216

D10 = 47

D7 + D8 + D9 = 216 - 47 = 169

D6 = (D6 + D7 + D8 + D9) - (D7 + D8 + D9 ) = 196 - 169 = 27